a farmer mixes wheat based feed (12% protein) with a corn based feed (7% protein) to make 610 pounds of a 10.5% protein feed. How many pounds of each type of feed does he use to make the mixture
Answer by richwmiller(17219)

a+b=610,
0.12*a+0.07*b=0.105*610
a=610-b
0.12*(610-b)+0.07*b=64.05
73.2-0.12b+0.07*b=64.05
-0.05*b=-9.15
b=183.0
a=610-b
a=427.0 lb at 12% b=183.0 lb at 7%
check
0.12*427.0+0.07*183.0=0.105*610
51.24+12.81=64.05
64.05=64.05
ok